OpenStudy (anonymous):

1. What is the sum of the geometric sequence 8, -16, 32 . if there are 15 terms? (1 point) 2. What is the sum of the geometric sequence 4, 12, 36 . if there are 9 terms? (1 point) 3. What is the sum of a 6-term geometric sequence if the first term is 11, the last term is -11,264 and the common ratio is -4? (1 point) 4. What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250?

OpenStudy (anonymous):

For the 1st geometric sequence, the common ratio is q=-2 -16/8 = -2 32/-16=-2 The sum of 15 terms could be found using the formula: S = 8*[(-2)^15-1]/(-2-1)

OpenStudy (anonymous):

For the 2nd geometric series, the ratio is q = 3 s = 4*(3^9 - 1)/(3-1) s = 2*(3^9 - 1)

OpenStudy (anonymous):

how do you figure that out? and whats the formula?

OpenStudy (anonymous):

Formula for calulating the sum of n terms of geometric sequence is: Sn = b1*(q^n - 1)/(q-1), where b1 is the 1st term and q is the common ratio

OpenStudy (anonymous):

For the 1st geometric series, the 1st term b1 is 8

OpenStudy (anonymous):

For the 2nd geom. series, the first term is 4

OpenStudy (anonymous):

okay, i think i understand, how would yuou solve 3 and 4 with a term?

OpenStudy (anonymous):

Well. we notice that to apply the formula we need to know only the 1st term and the ratio, we don't need anything else.

OpenStudy (anonymous):

S6 = 11*[(-4)^6-1]/(-4-1) S6 = (-11/5)*(4^6 - 1)

OpenStudy (anonymous):

For the 4th geom. series, we need to determine the common ratio. We know the last term and the first term and the fomrula for the general term; an = a1*q^(n-1) 781,250 = 10*q^(n-1) q^(n-1) = 78,125